# Training a Utility-based Retriever Through Shared Context Attribution for Retrieval-Augmented Language Models

### v1

@@ 4.2 Baselines
Retrievers are optimized using feedback from the generator. We select two reproducible methods, RePlug (Shi et al., 2023) and AAR (Yu et al., 2023), both of which are initialized from Contriever.
We utilize LLaMA-3-8B-Instruct (AI@Meta, 2024) and Qwen2.5-3B-Instruct (Team, 2024) as the generators in RALMs. All retrieval-based baselines use the top-3 passages. Given that some retrievers may not be tuned by instructions, the query format for Contriever and its baselines only contains $x$, without task instruction $I$. For BGE and its baselines, the query foramt follows the definition in Section 3.1, which contains both $x$ and $I$.
@@ 5 Results
In this section, we present the results of our experiment (§5.1), and conduct ablation study (§5.2), retrieval evaluation (§5.3), and case study (§5.4).
@@ 5.1 Overall Performance
In the evaluation on six in-domain datasets, the retrievers trained by SCARLet achieves the best performance in five datasets when using LLaMA-3-8B as the generator, and in four datasets when using Qwen-2.5-3B as the generator. Except for NQ and FEVER, SCARLet consistently outperforms the initial baselines, including Contriever and BGE.
@@ 5.2 Ablation Study
Table 5: Evaluation results on $\mathbb{X}^{2}$-Retrieval, averaged nDCG@10 for each datasets. Bracketed values indicate the changes in metrics compared to the initial model. The experiment follows the closed setup
Figure 5: Case Study on HotpotQA. The passage is from the corpus and has varying recall rankings for different retrievers. Orange text indicates necessary reasoning information for the question.
5.3 Aspects of Retrieval Utility
GTI
This benchmark was introduced in Section 3.4. Its primary goal is to evaluate whether the retriever can bypass pitfalls of semantic relevance and prioritize retrieving passages that are really useful for answering questions.
Bright
$\mathbb{X}^{2}$-Retrieval
@@ 5.4 Case Study
Multi-hop QA is a task that requires multiple pieces of information and multi-step reasoning to solve (Mavi et al., 2024). Given the characteristics of the task, we believe that retrieval utility should point to passages that may contain information necessary for the reasoning chain. We select a representative example from the test split of the HotpotQA dataset, as shown in Figure 5. To answer the question, the reasoning chain is: knowing information about William Preston, identifying the 1996 American historical drama he appeared in, finding information about that drama, and determining its writer. Directly relevant information about William Preston is relatively easy to define. However, the shown passage which corresponds to the final reasoning step, has a poor match with the question in terms of semantic relevance. And BGE ranks it 8th. After training by SCARLet, the passage achieves a higher ranking of 3rd. For more case studies, please refer to Appendix E.
@@ 6 Conclusion
This study focuses on utility-based retrieval, a paradigm that moves beyond semantic relevance to prioritize downstream task performance in RALMs. We highlight two key challenges faced by existing research. To solve the limitations, we propose SCARLet, a novel framework to enhance utility-based retrieval. To mitigate semantic interference on utility features during training, SCARLet incorporates a data synthesis pipeline, which narrows the semantic gap between different tasks. To address the issue of inaccurate passage-level utility estimation, SCARLet employs a perturbation-based attribution method to capture the synergy between passages in the context. Lastly, SCARLet utilizes a one-dimensional clustering method to sample positive and negative passages from the shared context for retriever optimization. Through experiments, we demonstrate that SCARLet can effectively enhance the overall task performance of RALMs, and can also bring improvements in complex retrieval benchmarks. We hope this study can inspires further research on utility-based retrieval.
@@
In this study, the scope of downstream tasks is limited, covering only several classic datasets. We believe that tasks should not be restricted to existing datasets, and maybe incorporating a task augmentation stage could further enhance generalization, which we leave for future work. In addition, due to memory and time constraints, this study does not evaluate larger-scale retrievers and generators. Furthermore, apart from GPT-4o, we only try GPT-4o-mini as the synthesizer, which performed poorly. Models with stronger reasoning capabilities may synthesize higher-quality data, potentially leading to greater performance improvements.
@@
The purpose of this study is to enhance the performance of RALMs in several common NLP tasks. All datasets and corpora involved are publicly available, and we ensure that all used data comply with the usage and privacy policies established by the original authors. The synthetic data in our method is exclusively used for training the retriever model. Moreover, given the security assurance of the synthesizer model, the probability of generating harmful passages and data is extremely minimal.
@@ Appendix C Details of Implementation
During the data synthesis stage, the temperature of the synthesizer model is set to 0.5. In the utility attribution stage, the number of sampled perturbation vectors $n$ is set to 64, with a perturbation probability of 0.5. During training, we set the learning rate as 6e-5, and epochs as 3. All experiments are conducted on NVIDIA A100 GPUs in torch.float32 precision.
@@ Appendix D Additional Experimental Results
The results presented in Table 2 are under the closed setup. In contrast, the pooled setup refers to merging the corpora of different datasets into a single corpus, where all retrieval is performed with the unified corpus. This setup better simulates real-world retrieval scenarios and enables a fairer evaluation of generalization. The results under the pooled setup are shown in Table 13. All baselines perform similarly to those in the closed setup, and some outperform them, demonstrating generalization on the unified corpus.
@@ Appendix F Example of Shared Context
Table 13: Results of the main experiment in the pooled setup.

### v2

@@ 4.2 Baselines
Retrievers are optimized using feedback from the generator. We select two recent methods, RePlug (Shi et al., 2023) and AAR (Yu et al., 2023), both of which are initialized from Contriever.
We utilize LLaMA-3-8B-Instruct (AI@Meta, 2024) and Qwen2.5-3B-Instruct (Team, 2024) as the generators in RALMs. All retrieval-based baselines use the top-3 passages. Given that some retrievers may not be tuned by instructions, the query format for Contriever and its baselines only contains $x$, without task instruction $I$. For BGE and its baselines, the query format follows the definition in Section 3.1, which contains both $x$ and $I$.
@@ 5 Results
In this section, we present the results of main experiment (§5.1), ablation study (§5.2), retrieval evaluation (§5.3), and case study (§5.4).
@@ 5.1 Overall Performance
In the evaluation on six in-domain datasets, the retrievers trained by SCARLet achieve the best performance in five datasets when using LLaMA-3-8B as the generator, and in four datasets when using Qwen-2.5-3B as the generator. Except for NQ and FEVER, SCARLet consistently outperforms the initial baselines, including Contriever and BGE.
@@ 5.2 Ablation Study
5.3 Aspects of Retrieval Utility
GTI
This benchmark was introduced in Section 3.4. Its goal is to evaluate whether retrievers can bypass pitfalls of semantic relevance and prioritize passages that are useful for answering questions.
Bright
$\mathbb{X}^{2}$-Retrieval
@@ 5.3 Aspects of Retrieval Utility
Contriever 10.5 19.6 6.9
$\textrm{SCARLet}_{\textrm{Contriever}}$ $\textrm{13.3}_{{\color[rgb]{0.234375,0.51171875,0.01953125}\left(+2.8\right)}}$ $\textrm{19.2}_{{\color[rgb]{0.5,0.01171875,0.1796875}\left(-0.4\right)}}$ $\textrm{8.7}_{{\color[rgb]{0.234375,0.51171875,0.01953125}\left(+1.8\right)}}$
@@
Table 5: Evaluation results on $\mathbb{X}^{2}$-Retrieval, averaged nDCG@10 for each datasets. Bracketed values indicate the changes in metrics compared to the initial model.
@@ 5.4 Case Study
Multi-hop QA is a task that requires multiple pieces of information and multi-step reasoning to derive the solution (Mavi et al., 2024). Given the characteristics of the task, we believe that retrieval utility should point to passages that may contain information necessary for the reasoning chain. We select a representative example from the test split of the HotpotQA dataset, as shown in Figure 5. To answer the question, the reasoning chain is: knowing information about William Preston, identifying the 1996 American historical drama he appeared in, finding information about that drama, and determining its writer. Directly relevant information about William Preston is relatively easy to define. However, the shown passage which corresponds to the final reasoning step, has a poor match with the question in terms of semantic relevance. And BGE ranks it 8th. After training by SCARLet, the passage achieves a higher ranking of 3rd. For more case studies, please refer to Appendix E.
Figure 5: Case Study on HotpotQA. The passage is ranked variously by different retrievers. Orange text indicates necessary reasoning information.
@@ 6 Conclusion
This study focuses on utility-based retrieval, a paradigm that moves beyond semantic relevance to prioritize downstream task performance in RALMs. We highlight two key challenges faced by existing research. To solve the limitations, we propose SCARLet, a novel framework to enhance utility-based retrieval. To mitigate semantic interference on utility features during training, SCARLet incorporates a shared context synthesis method, which narrows the semantic gap between different tasks. To address the issue of inaccurate passage-level utility estimation, SCARLet employs a perturbation-based attribution method to capture the synergy between passages. Lastly, SCARLet utilizes a one-dimensional clustering method to sample positive and negative passages for retriever optimization. Through experiments, we demonstrate that SCARLet can effectively enhance the overall performance of RALMs, and brings improvements in complex retrieval benchmarks. We hope this study can inspire further research on utility-based retrieval.
@@
This study only covers several classic downstream datasets. We believe that incorporating a task augmentation stage could further enhance generalization, which we leave for future work. Moreover, there is a noticeable decline in retrieval performance in the code domain during generalization tests. Therefore, future work should also focus on improving the integration of different corpus structures. In addition, due to environmental constraints, this study does not evaluate larger-scale retrievers and generators. Furthermore, we also try GPT-4o-mini as the synthesizer, which performed poorly. Thus our framework should be equipped with models with stronger reasoning capabilities.
@@
All datasets and corpora involved in this study are publicly available, and we ensure that all used data comply with the usage and privacy policies established by the original authors. The synthetic data is exclusively used for training the retriever model. Moreover, given the security assurance of the synthesizer model, the probability of generating harmful passages and data is extremely minimal.
Acknowledgments
This work was supported by the National Key R&D Program of China (2023YFC3303800).
@@ Appendix C Details of Implementation
During the data synthesis stage, the temperature of the synthesizer model is set to 0.5. In the utility attribution stage, the number of sampled perturbation vectors $n$ is set to 64, with a perturbation probability of 0.5. During training, we set the learning rate as 6e-5, and epochs as 1. All experiments are conducted on NVIDIA A100 GPUs in torch.float32 precision.
@@ Appendix D Additional Experimental Results
The results presented in Table 2 are under the closed corpus setup, i.e., retrievers search passages only from the corpus of the corresponding dataset. In contrast, the pooled corpus setup refers to merging the corpora of different datasets into a single corpus, where all retrieval is performed with the unified corpus. This setup better simulates real-world retrieval scenarios and enables a fairer evaluation of generalization. The experimental results under the pooled corpus setup are shown in Table 13. All baselines perform similarly to those in the closed corpus setup, and some outperform them, demonstrating generalization of SCARLet on the unified corpus.
@@ Appendix F Example of Shared Context
Table 13: Results of the main experiment in the pooled corpus setup. The unified corpus includes corpora of Wikipedia dump, BeIR-SciFact, BeIR-ClimateFEVER and BeIR-FiQA.